# Shuchang Zhang, Hui Zhang, Hongxia Wang

### v4

@@ 3.2 Convergence rate
where $\gamma\in(0,1]$, the sequence $\{b_{k}\}_{k\in\mathbb{N}}$ is defined by $b_{k}=\min\{\frac{2}{\gamma k},1\}$, and $\{c_{k}\}_{k\in\mathbb{N}}$ is a sequence of real numbers such that $c_{k}\leq M_{1}<\infty$. Then the sequence $\{a_{k}\}_{k\in\mathbb{N}}$ satisfies
@@
By applying the Lemma 3.9, we will establish a sublinear convergence rate for $\left\|\mathbf{u}^{k}-\mathbf{u}^{k-1}\right\|$ and fixed point residual $\left\|\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\right\|$ or $\left\|\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\right\|_{\mathcal{M}}$.
@@
Let $\mathcal{A}:\mathcal{H}\to 2^{\mathcal{H}}$ be a maximal operator with $\mathrm{zer}\mathcal{A}\neq\emptyset$, and $\mathcal{M}$ an admissible preconditioner such that $\mathcal{M}^{-1}\mathcal{A}$ is $\mathcal{M}$-monotone and $(\mathcal{M}+\mathcal{A})^{-1}$ is $L$-Lipshitz. Let $\{u^{k}\}$ be the sequence generated by (5). If $\mu_{k}=\min\{\frac{2}{k},1\}$, then
@@
    $\left\|\mathbf{u}^{k}-\mathbf{u}^{k-1}\right\|_{\mathcal{M}}\leq\frac{2M}{k}$, i.e., $\left\|\mathbf{u}^{k}-\mathbf{u}^{k-1}\right\|_{\mathcal{M}}=\mathcal{O}(\frac{1}{k})$.
@@
    $\left\|\mathbf{u}^{k+1}-\mathbf{u}^{k}\right\|=\mathcal{O}(\frac{1}{k})$, $\left\|\mathbf{u}^{k}-\mathcal{T}\mathbf{u}^{k}\right\|=\mathcal{O}(\frac{1}{k})$.
@@
where $M=\left\|\mathbf{a}-\mathbf{u}^{*}\right\|_{\mathcal{M}}+C_{1},\mathbf{u}^{*}\in\mathrm{Fix}(\mathcal{T})$, and $C_{1}$ is defined by (27). Applying Lemma 3.9 together with $c_{k}=M_{1}=M,J=2,\gamma=1,a_{k}=\left\|\mathbf{u}^{k}-\mathbf{u}^{k-1}\right\|_{\mathcal{M}}$ and $\mu_{k}=b_{k}$ to the above inequality, we have $\left\|\mathbf{u}^{k}-\mathbf{u}^{k-1}\right\|_{\mathcal{M}}\leq\frac{2M}{k}$.
@@
$\leq\frac{2MC}{k}+\frac{2M^{\prime}}{k(k+1)}=\mathcal{O}(\frac{1}{k}),$
@@
$\leq\frac{2MC}{k}+\frac{2M^{\prime}}{k(k+1)}+\frac{2M^{\prime}}{k+1}=\mathcal{O}(\frac{1}{k}).$
Let $\mathcal{M}=\mathcal{C}\mathcal{C}^{*}$ be a decomposition of $\mathcal{M}$ according to [9, Proposition 2.3]. By [9, Theorem 2.13], the operator $\tilde{\mathcal{T}}=\mathcal{C}^{*}(\mathcal{M}+\mathcal{A})^{-1}\mathcal{C}$ is FNE, let $\mathbf{w}^{k}=\mathcal{C}^{*}\mathbf{u}^{k}$ and $\mathbf{a}=\mathbf{u}^{0}$, then HPPP is equivalent to the following reduced algorithm:
@@
Let $\mathcal{A}:\mathcal{H}\to 2^{\mathcal{H}}$ be a maximal operator with $\mathrm{zer}\mathcal{A}\neq\emptyset$, and $\mathcal{M}$ an admissible preconditioner such that $\mathcal{M}^{-1}\mathcal{A}$ is $\mathcal{M}$-monotone and $(\mathcal{M}+\mathcal{A})^{-1}$ is $L$-Lipshitz. Let $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}},\{\mathbf{w}^{k}\}_{k\in\mathbb{N}}$ be the sequences generated by (5) and (16). If $\mu_{k}=\frac{1}{k+1}$ and $\mathbf{a}=\mathbf{u}^{0}$, then
@@
    $\left\|\mathbf{w}^{k}-\tilde{\mathcal{T}}\mathbf{w}^{k}\right\|\leq\frac{2\left\|\mathbf{w}^{0}-\mathbf{w}^{*}\right\|}{k+1}$ for $k\geq 0$ and $\mathbf{w}^{*}\in\mathrm{Fix}(\tilde{\mathcal{T}})$.
@@ 3.3 GraRED-HP3
The primal-dual algorithm to solve (4) is viewed as the fixed-point iteration $\mathbf{u}^{k+1}=\mathcal{T}\mathbf{u}^{k}=(\mathcal{A}+\mathcal{M})^{-1}\mathcal{M}\mathbf{u}^{k}$ with $\mathcal{A}=\begin{pmatrix}\lambda\partial f&\mathbf{K}^{*}\\ -\mathbf{K}&\partial g^{*}\end{pmatrix},\mathcal{M}=\begin{pmatrix}\frac{1}{\tau}I&-\mathbf{K}^{*}\\ -\mathbf{K}&\frac{1}{s}I\end{pmatrix}$. Under the degenerate case $\tau s\left\|\mathbf{K}\right\|^{2}=1$, the HPPP iteration is given by
@@
where $\mathbf{a}=(\mathbf{x}_{a},\mathbf{y}_{a}),(\mathbf{x}_{0},\mathbf{y}_{0})\in\mathcal{X}\times\mathcal{Y}$ are the anchor and initial points. Under the ideal conditions of RED and with Assumption (A), according to Lemma 2.6, both the denoiser $D_{\sigma}$ and the residual $R$ can be expressed as the proximal operators of implicit regularizers, i.e., $D_{\sigma}(\mathbf{x})=\mathrm{prox}_{\phi}(\mathbf{x}),R(\mathbf{x})=\mathrm{prox}_{\phi^{*}}(\mathbf{x})$. Let $g(\mathbf{x})=\frac{\phi(s\mathbf{x})}{s}$, then $\phi^{*}=sg^{*}$ [5, Theorem 4.14], we can replace $(I+s\partial g^{*})^{-1}$ with the residual $R$ in (17). Set $\mathbf{K}=I$, the implicit Gradient RED via HPPP called GraRED-HP3 is proposed in Algorithm 3. Even if the conditions of RED are difficult to meet [45], recent developments [49, 54, 42, 21, 15, 10, 57] can learn FNE denoisers. According to Lemma 2.8, both the denoiser and the residual still act as the proximal operators of implicit regularizers.
@@ 3.3.1 Acceleration for PnP methods
PnP-ADMM (see Algorithm 1) is a well-known method to solve (1), which can be written into an equivalent DRS form [49, 33], i.e., $\mathbf{w}^{k+1}=\tilde{\mathcal{T}}\mathbf{w}^{k}=\mathbf{w}^{k}+D_{\sigma}(2\mathrm{prox}_{\lambda f}(\mathbf{w}^{k})-\mathbf{w}^{k})-\mathrm{prox}_{\lambda f}(\mathbf{w}^{k})$. When $\mathbf{w}^{k}=\mathbf{x}^{k}-\mathbf{y}^{k},\tau=s=1,\mu_{k}=0$ in GraRED-HP3 results exactly in the DRS iteration. Thus, PnP-DRS is a special case of GraRED-HP3, which can be obtained from the perspective of PPP [11, 9]. The related GraRED-P3 algorithm is shown in Algorithm 4. Furthermore, Proposition 3.12 demonstrates that GraRED-HP3 is an accelerated PnP method achieving an $\mathcal{O}(1/k)$ convergence rate for the fixed-point residual.
@@
In [28], a worst-case convergence rate of $\mathcal{O}(\frac{1}{\sqrt{k}})$ for $\left\|\mathbf{w}^{k}-\tilde{\mathcal{T}}\mathbf{w}^{k}\right\|$ was established for DRS, which also applies to PnP-ADMM. In contrast, GraRED-HP3 obtain an accelerated convergence rate of $\mathcal{O}(\frac{1}{k})$ for $\left\|\mathbf{w}^{k}-\tilde{\mathcal{T}}\mathbf{w}^{k}\right\|$.
@@
1:Input: $\mathbf{x}^{0}\in\mathbb{R}^{n}$, $\{t_{k}\}_{k\in\mathbb{N}}$, $\alpha\in(0,1)$, $\lambda$, $\mu$, $J$, $N,\delta>0$ and the denoiser $T$.
@@ 3.3.2 Compared with RED-PRO
RED-PRO sheds new insights on denoiser prior, which theoretically bridges between RED and PnP prior. However, practical denoisers often have narrow fixed-point sets, leading to suboptimal recovery solutions. To address this issue, they relax the hard constraint of $\mathrm{Fix}(T)$ and replace it with a dilated fixed-point set, defined for some $\delta>0$ as
@@
Algorithm 5 outlines RRP, where $T_{\alpha}=\alpha T+(1-\alpha)I$ with the demi-contractive denoiser $T$. When the steepest descent method solves (19), it requires an inner loop to calculate the fixed-point projection $P_{\mathrm{Fix}(T)}(\mathbf{x}^{k})$, and the line 5 in RRP is equivalent to the Halpern iteration.
Comparing Algorithm 3 with the inner loop of Algorithm 5, both methods share the same algorithmic form of Halpern iteration and utilize identical anchor coefficients. In both cases, finding $P_{\mathrm{Fix}(T)}(\mathbf{x}^{k})$ or $P_{\mathrm{Fix}(\mathcal{T})}^{\mathcal{M}}(\mathbf{a})$ corresponds to a bilevel optimization problem that finds the point in the fixed-point set closest to the anchor point $\mathbf{x}^{k}$ or $\mathbf{a}$.
The key difference is that HPPP extends Halpern iteration to a degenerate form by employing a positive semi-definite metric $\mathcal{M}$, whereas the denoiser operator $T$ in RED-PRO is defined under the standard metric $\mathcal{M}=I$. From an optimization perspective, the RRP algorithm is based on gradient descent and uses the classic Halpern iteration in its inner loop. In contrast, $\mathcal{T}$ in GraRED-HP3 is an algorithmic operator and has no inner loop. From a PPP standpoint, we integrate HPPP with denoiser priors to propose Algorithm 3.
@@ 4 Experiments
In this section, we show the numerical experiments of the algorithms discussed in section 3. Firstly, we will verify the implicit regularity of HPPP by an easy 1D example, and validate the accelerated advantage of HPPP and Restarted HPPP by other toy examples. Then, we compare the CP, PPP (8), GraRED-HP3, and GraRED-P3 for image deblurring and inpainting under the same setting, and verify the efficiency of the proposed algorithms. The source code is available at https://github.com/zsc15/HPPP.
@@ 4.1 A toy example
We consider the following optimization problem in $\mathbb{R}$, i.e.,
@@
where $f(x)=\max\{-x,0\}$ and $g(x)=\max\{1-x,0\}$, we plot $f(x)+g(x)$ in Figure 1.
@@
The corresponding saddle-point problem is $\min_{x\in\mathbb{R}}\max_{y\in\mathbb{R}}\{xy+f(x)-g^{*}(y)\}$, where $g^{*}(y)=\max_{x}[yx-\max\{1-x,0\}]=y+\delta_{[-1,0]}(y)$ and $\delta_{[-1,0]}$ is the indicator function of the interval $[-1,0]$. We denote the optimal set $X^{*}=[1,+\infty)=\arg\min_{x\mathbb{\in\mathbb{R}}}f(x)+g(x)$ and the primal-dual objective function $F(x,y)=xy+f(x)-g^{*}(y)$. Let us solve the saddle-point set $\{(x^{*},y^{*}):F(x^{*},y)\leq F(x^{*},y^{*})\leq F(x,y^{*}),\forall(x,y)\in\mathbb{R}^{2}\}$ of $F(x,y)$. Fixed $x^{*}\geq 1$, then
@@
If $x^{*}=1$, assume that $-1\leq y^{*}<0$, then $F(1,y^{*})=0$, while $F(x,y^{*})=y^{*}(x-1)+\max\{-x,0\}$ and there exists $x=2$ such that $F(2,y^{*})=y^{*}<F(1,y^{*})$, which leads to contradiction. Therefore, the saddle-point set is $\Omega=\{(x^{*},y^{*}):x^{*}\geq 1,y^{*}=0\}$.
@@
Figure 2: $\mathcal{M}$-projection onto the saddle point set.

### v5

@@ 3.2 Convergence rate
where $\gamma\in(0,1]$, the sequence $\{b_{k}\}_{k\in\mathbb{N}}$ is defined as in $b_{k}=\min\{\frac{2}{\gamma k},1\}$, and $\{c_{k}\}_{k\in\mathbb{N}}$ is a sequence of real numbers such that $c_{k}\leq M_{1}<\infty$. Then the sequence $\{a_{k}\}_{k\in\mathbb{N}}$ satisfies
@@
By applying Lemma 3.9, we will establish a sublinear convergence rate for $\left\|\mathbf{u}^{k}-\mathbf{u}^{k-1}\right\|$ and fixed-point residual $\left\|\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\right\|$ or $\left\|\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\right\|_{\mathcal{M}}$.
@@
Let $\mathcal{A}:\mathcal{H}\to 2^{\mathcal{H}}$ be a maximal operator with $\mathrm{zer}\mathcal{A}\neq\emptyset$, and let $\mathcal{M}$ be an admissible preconditioner such that $\mathcal{M}^{-1}\mathcal{A}$ is $\mathcal{M}$-monotone and $(\mathcal{M}+\mathcal{A})^{-1}$ is $L$-Lipschitz. Let $\{u^{k}\}$ be the sequence generated by (5). If $\mu_{k}=\min\{\frac{2}{k},1\}$, then
@@
    $\left\|\mathbf{u}^{k}-\mathbf{u}^{k-1}\right\|_{\mathcal{M}}\leq\frac{2M}{k}$, i.e., $\left\|\mathbf{u}^{k}-\mathbf{u}^{k-1}\right\|_{\mathcal{M}}=\mathcal{O}\left(\frac{1}{k}\right)$;
@@
    $\left\|\mathbf{u}^{k+1}-\mathbf{u}^{k}\right\|=\mathcal{O}\left(\frac{1}{k}\right)$, $\left\|\mathbf{u}^{k}-\mathcal{T}\mathbf{u}^{k}\right\|=\mathcal{O}\left(\frac{1}{k}\right)$.
@@
where $M=\left\|\mathbf{a}-\mathbf{u}^{*}\right\|_{\mathcal{M}}+C_{1},\mathbf{u}^{*}\in\mathrm{Fix}(\mathcal{T})$, and $C_{1}$ is defined as in (27). Applying Lemma 3.9 together with $c_{k}=M_{1}=M,J=2,\gamma=1,a_{k}=\left\|\mathbf{u}^{k}-\mathbf{u}^{k-1}\right\|_{\mathcal{M}}$, and $\mu_{k}=b_{k}$ to the above inequality, we have $\left\|\mathbf{u}^{k}-\mathbf{u}^{k-1}\right\|_{\mathcal{M}}\leq\frac{2M}{k}$.
@@
$\leq\frac{2MC}{k}+\frac{2M^{\prime}}{k(k+1)}=\mathcal{O}\left(\frac{1}{k}\right),$
@@
$\leq\frac{2MC}{k}+\frac{2M^{\prime}}{k(k+1)}+\frac{2M^{\prime}}{k+1}=\mathcal{O}\left(\frac{1}{k}\right).$
Let $\mathcal{M}=\mathcal{C}\mathcal{C}^{*}$ be a decomposition of $\mathcal{M}$ according to [9, Proposition 2.3]. By [9, Theorem 2.13], the operator $\tilde{\mathcal{T}}=\mathcal{C}^{*}(\mathcal{M}+\mathcal{A})^{-1}\mathcal{C}$ is FNE. Let $\mathbf{w}^{k}=\mathcal{C}^{*}\mathbf{u}^{k}$ and $\mathbf{a}=\mathbf{u}^{0}$, and then HPPP is equivalent to the following reduced algorithm:
@@
Let $\mathcal{A}:\mathcal{H}\to 2^{\mathcal{H}}$ be a maximal operator with $\mathrm{zer}\mathcal{A}\neq\emptyset$, and let $\mathcal{M}$ be an admissible preconditioner such that $\mathcal{M}^{-1}\mathcal{A}$ is $\mathcal{M}$-monotone and $(\mathcal{M}+\mathcal{A})^{-1}$ is $L$-Lipschitz. Let $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}},\{\mathbf{w}^{k}\}_{k\in\mathbb{N}}$ be the sequences generated by (5) and (16). If $\mu_{k}=\frac{1}{k+1}$ and $\mathbf{a}=\mathbf{u}^{0}$, then
@@
    $\left\|\mathbf{w}^{k}-\tilde{\mathcal{T}}\mathbf{w}^{k}\right\|\leq\frac{2\left\|\mathbf{w}^{0}-\mathbf{w}^{*}\right\|}{k+1}$ for $k\geq 0$ and $\mathbf{w}^{*}\in\mathrm{Fix}(\tilde{\mathcal{T}})$,
@@ 3.3 GraRED-HP3
The primal-dual algorithm for solving (4) is viewed as the fixed-point iteration $\mathbf{u}^{k+1}=\mathcal{T}\mathbf{u}^{k}=(\mathcal{A}+\mathcal{M})^{-1}\mathcal{M}\mathbf{u}^{k}$ with $\mathcal{A}=\begin{pmatrix}\lambda\partial f&\mathbf{K}^{*}\\ -\mathbf{K}&\partial g^{*}\end{pmatrix},\mathcal{M}=\begin{pmatrix}\frac{1}{\tau}I&-\mathbf{K}^{*}\\ -\mathbf{K}&\frac{1}{s}I\end{pmatrix}$. Under the degenerate case $\tau s\left\|\mathbf{K}\right\|^{2}=1$, the HPPP iteration is given by
@@
where $\mathbf{a}=(\mathbf{x}_{a},\mathbf{y}_{a}),(\mathbf{x}_{0},\mathbf{y}_{0})\in\mathcal{X}\times\mathcal{Y}$ are the anchor and initial points. Under the ideal conditions of RED and with assumption (A), according to Lemma 2.6, both the denoiser $D_{\sigma}$ and the residual $R$ can be expressed as the proximal operators of implicit regularizers, i.e., $D_{\sigma}(\mathbf{x})=\mathrm{prox}_{\phi}(\mathbf{x}),R(\mathbf{x})=\mathrm{prox}_{\phi^{*}}(\mathbf{x})$. Let $g(\mathbf{x})=\frac{\phi(s\mathbf{x})}{s}$; then $\phi^{*}=sg^{*}$ [5, Theorem 4.14], and we can replace $(I+s\partial g^{*})^{-1}$ with the residual $R$ in (17). Set $\mathbf{K}=I$; the implicit gradient RED via HPPP called GraRED-HP3 is proposed in Algorithm 3. Even if the conditions of RED are difficult to meet [45], recent developments [49, 54, 42, 21, 15, 10, 57] can learn FNE denoisers. According to Lemma 2.8, both the denoiser and the residual still act as the proximal operators of implicit regularizers.
@@ 3.3.1 Acceleration for PnP methods
PnP-ADMM (see Algorithm 1) is a well-known method for solving (1), which can be written into an equivalent DRS form [49, 33], i.e., $\mathbf{w}^{k+1}=\tilde{\mathcal{T}}\mathbf{w}^{k}=\mathbf{w}^{k}+D_{\sigma}(2\mathrm{prox}_{\lambda f}(\mathbf{w}^{k})-\mathbf{w}^{k})-\mathrm{prox}_{\lambda f}(\mathbf{w}^{k})$. When $\mathbf{w}^{k}=\mathbf{x}^{k}-\mathbf{y}^{k},\tau=s=1,\mu_{k}=0$ in GraRED-HP3 results exactly in the DRS iteration. Thus, PnP-DRS is a special case of GraRED-HP3, which can be obtained from the perspective of PPP [11, 9]. The related GraRED-P3 algorithm is shown in Algorithm 4. Furthermore, Proposition 3.12 demonstrates that GraRED-HP3 is an accelerated PnP method achieving an $\mathcal{O}(1/k)$ convergence rate for the fixed-point residual.
@@
In [28], a worst-case convergence rate of $\mathcal{O}\left(\frac{1}{\sqrt{k}}\right)$ for $\left\|\mathbf{w}^{k}-\tilde{\mathcal{T}}\mathbf{w}^{k}\right\|$ was established for DRS, which also applies to PnP-ADMM. In contrast, GraRED-HP3 obtains an accelerated convergence rate of $\mathcal{O}\left(\frac{1}{k}\right)$ for $\left\|\mathbf{w}^{k}-\tilde{\mathcal{T}}\mathbf{w}^{k}\right\|$.
@@
1:Input: $\mathbf{x}^{0}\in\mathbb{R}^{n}$, $\{t_{k}\}_{k\in\mathbb{N}}$, $\alpha\in(0,1)$, $\lambda$, $\mu$, $J$, $N,\delta>0$, and the denoiser $T$.
@@ 3.3.2 Compared with RED-PRO
RED-PRO sheds new insights on the denoiser prior, which theoretically bridges the gap between RED and the PnP prior. However, practical denoisers often have narrow fixed-point sets, leading to suboptimal recovery solutions. To address this issue, they relax the hard constraint of $\mathrm{Fix}(T)$ and replace it with a dilated fixed-point set, defined for some $\delta>0$ as
@@
Algorithm 5 outlines RRP, where $T_{\alpha}=\alpha T+(1-\alpha)I$ with the demicontractive denoiser $T$. When the steepest descent method solves (19), it requires an inner loop to calculate the fixed-point projection $P_{\mathrm{Fix}(T)}(\mathbf{x}^{k})$, and line 5 in RRP is equivalent to the Halpern iteration.
Comparing Algorithm 3 with the inner loop of Algorithm 5, we see that both methods share the same algorithmic form of Halpern iteration and utilize identical anchor coefficients. In both cases, finding $P_{\mathrm{Fix}(T)}(\mathbf{x}^{k})$ or $P_{\mathrm{Fix}(\mathcal{T})}^{\mathcal{M}}(\mathbf{a})$ corresponds to a bilevel optimization problem that finds the point in the fixed-point set closest to the anchor point $\mathbf{x}^{k}$ or $\mathbf{a}$.
The key difference is that HPPP extends Halpern iteration to a degenerate form by employing a positive semidefinite metric $\mathcal{M}$, whereas the denoiser operator $T$ in RED-PRO is defined under the standard metric $\mathcal{M}=I$. From an optimization perspective, the RRP algorithm is based on gradient descent and uses the classic Halpern iteration in its inner loop. In contrast, $\mathcal{T}$ in GraRED-HP3 is an algorithmic operator and has no inner loop. From a PPP standpoint, we integrate HPPP with denoiser priors to propose Algorithm 3.
@@ 4 Experiments
In this section, we show the numerical experiments of the algorithms discussed in section 3. First, we will verify the implicit regularity of HPPP by an easy 1D example, and validate the accelerated advantage of HPPP and restarted HPPP by other toy examples. Then, we compare CP, PPP (8), GraRED-HP3, and GraRED-P3 for image deblurring and inpainting under the same setting and verify the efficiency of the proposed algorithms. The source code is available at https://github.com/zsc15/HPPP.
@@ 4.1 A toy example
We consider the optimization problem in $\mathbb{R}$, i.e.,
@@
where $f(x)=\max\{-x,0\}$ and $g(x)=\max\{1-x,0\}$, and we plot $f(x)+g(x)$ in Figure 1.
@@
The corresponding saddle-point problem is $\min_{x\in\mathbb{R}}\max_{y\in\mathbb{R}}\{xy+f(x)-g^{*}(y)\}$, where $g^{*}(y)=\max_{x}[yx-\max\{1-x,0\}]=y+\delta_{[-1,0]}(y)$ and $\delta_{[-1,0]}$ is the indicator function of the interval $[-1,0]$. We denote the optimal set $X^{*}=[1,+\infty)=\arg\min_{x\mathbb{\in\mathbb{R}}}f(x)+g(x)$ and the primal-dual objective function $F(x,y)=xy+f(x)-g^{*}(y)$. Let us solve the saddle-point set $\{(x^{*},y^{*}):F(x^{*},y)\leq F(x^{*},y^{*})\leq F(x,y^{*})\text{ for all }(x,y)\in\mathbb{R}^{2}\}$ of $F(x,y)$. Fix $x^{*}\geq 1$; then
@@
If $x^{*}=1$, assume that $-1\leq y^{*}<0$, then $F(1,y^{*})=0$, while $F(x,y^{*})=y^{*}(x-1)+\max\{-x,0\}$, and there exists $x=2$ such that $F(2,y^{*})=y^{*}<F(1,y^{*})$, which leads to a contradiction. Therefore, the saddle-point set is $\Omega=\{(x^{*},y^{*}):x^{*}\geq 1,y^{*}=0\}$.
@@
Figure 2: $\mathcal{M}$-projection onto the saddle-point set.